# The Effect of Sampling Temperature on Problem Solving in Large Language Models

### v1

@@ Abstract
In this research study, we empirically investigate the effect of sampling temperature on the performance of Large Language Models (LLMs) on various problem-solving tasks. We created a multiple-choice question-and-answer (MCQA) exam by randomly sampling problems from standard LLM benchmarks. Then, we used four popular LLMs with five prompt-engineering techniques to solve the MCQA problems while increasing the sampling temperature from 0.0 to 1.0. Despite anecdotal reports to the contrary, our empirical results indicate that changes in temperature in the range 0.0 to 1.0 do not have a statistically significant impact on LLM performance for problem-solving tasks. In addition, these results appear to hold regardless of the LLM, the prompt-engineering technique, or the problem domain. All code, data, and supplemental materials are available on GitHub at: https://github.com/matthewrenze/jhu-llm-temperature.
Large Language Model, LLM, Generative Pretrained Transformer, GPT, sampling temperature
@@ 1.1 Background
In recent years, Large Language Models (LLMs) have revolutionized the field of artificial intelligence. Open-source LLMs and pay-per-use APIs have allowed engineers to incorporate LLMs in their AI systems. However, prompt engineering and hyperparameter tuning are required to work effectively with LLMs.
Prompt-engineering techniques help LLMs solve complex problems, avoid hallucinations, and provide more accurate responses. For example, we can use techniques like chain-of-thought, tree-of-thought, self-criticism, and self-consistency to improve LLM performance (Mialon et al., 2023; White et al., 2023).
In addition, several inference hyperparameters can be adjusted to change the LLM’s output at runtime. For example, hyperparameters like sampling temperature, top-k sampling, repetition penalty, and maximum token length all affect the LLMs output and performance (OpenAI, 2023a; Touvron et al., 2023; Wang et al., 2023a).
Despite significant interest in LLMs and progress in LLM best practices, many open questions remain about optimal prompt-engineering techniques and inference hyperparameters for LLMs. To complicate matters, various local optima may exist for each problem type and problem domain (Wang et al., 2023a).
@@ 1.2 Sampling Temperature
Sampling temperature is a hyperparameter of an LLM used in a temperature-based sampling process. It controls the randomness of the model’s output at inference time (Ackley et al., 1985; Hinton et al., 2015; Wang et al., 2020, 2023a).
At each step of an LLM’s decoding process, the LLM uses the previous tokens to choose the next output token. The final layer of the LLM uses a softmax function to convert raw scores (logits) into probabilities.
@@
A lower temperature makes the output of the LLM more deterministic, thus favoring the most likely predictions. This conservativeness is captured by the model’s tendency to produce more repetitive, focused, and less diverse output based on the patterns most commonly seen in the training data (Hinton et al., 2015; Wang et al., 2020, 2023a).
A higher temperature makes the output more random, thus favoring more “creative” predictions. This creativity is captured by the model’s willingness to explore more unconventional and less likely outputs. Higher temperatures can lead to novel text, diverse ideas, and creative solutions to problems (Hinton et al., 2015; Wang et al., 2020, 2023a).
@@ 1.3 Choosing a Sampling Temperature
Within the prompt-engineering community, there are a variety of opinions and best practices regarding the ideal sampling temperature for various problem-solving tasks (Microsoft, ; Shieh, ).
Low sampling temperatures are recommended for tasks requiring precision and factual accuracy, such as technical writing, code generation, or question-answering (Xu et al., 2022; Zhu et al., 2023). However, higher temperatures are recommended for tasks requiring creativity, such as writing poetry, creating stories, or brainstorming.
Higher temperatures also increase the probability of model hallucination. Hallucination is a phenomenon where an LLM produces statistically probable responses that are factually incorrect or nonsensical. As a result, optimal temperature selection is also a balance between creativity and hallucination (Lee, 2023).
Practical guidelines for choosing a sampling temperature for a specific task or problem domain are often vague or anecdotal. Prompt engineering guides often provide hypothetical examples of optimal sample temperatures for various tasks. However, they rarely cite any sources or provide empirical evidence [^1] A few empirical studies exist that indicate sampling temperature does have an effect on LLM performance on some types of problem-solving tasks (e.g., code generation, engineering exams, etc.) (Xu et al., 2022; Pursnani et al., 2023; Zhu et al., 2023).
As a result, the current state of choosing the optimal sample temperature for specific problems is largely based on guesswork, gut instinct, non-systematic experimentation, and iterative refinement [^2] Even the GPT-4 Technical Report explains that the authors used their “best-guess” when choosing sampling temperatures while evaluating GPT-4 on various benchmarks. See Appendix A in the GPT-4 Technical Report (OpenAI, 2023c)..
1.4 Hypothesis
Given the lack of strong evidence for best practices regarding optimal sampling temperature, our research study empirically investigated the effect of LLM sampling temperature on problem-solving tasks. To answer our primary research question, we formulated the following hypotheses:
Null Hypothesis ($H_{0}$): Increasing or decreasing sampling temperature in the range 0.0 to 1.0 has no effect on the problem-solving performance of the LLM.
Alternative Hypothesis ($H_{1}$): Increasing or decreasing sampling temperature in the range 0.0 to 1.0 improves the problem-solving performance of the LLM agent.
To test these hypotheses, we measured the LLM’s accuracy in correctly answering 1,000 multiple-choice questions across ten problem domains, ten times each. We then conducted a hypothesis test using a significance level of 0.05.
@@ 2.1 Data
The test dataset used in this research study consists of a series of Multiple-Choice Question-and-Answer (MCQA) exams derived from commonly used LLM performance benchmarks.
@@
Next, we selected a set of problems that covered a range of problem domains (e.g., math, science, law, etc.) In addition, we selected problem sets that spanned a range of difficulty levels (e.g., secondary school, university, etc.) These problem sets can be seen in Table 1.
@@
Table 1: Source of problem sets used to create the multi-domain MCQA exam.
Problem Set Benchmark Domain Questions Source
ARC Challenge Test ARC Science 1,173 (Clark et al., 2018)
AQUA-RAT AGI Eval Math 254 (Zhong et al., 2023)
Hellaswag Val Hellaswag Common Sense Reasoning 10,042 (Zellers et al., 2019)
LogiQA (English) AGI Eval Logic 651 (Zhong et al., 2023; Liu et al., 2020)
LSAT-AR AGI Eval Law (Analytic Reasoning) 230 (Zhong et al., 2023; Wang et al., 2021)
LSAT-LR AGI Eval Law (Logical Reasoning) 510 (Zhong et al., 2023; Wang et al., 2021)
LSAT-RC AGI Eval Law (Reading Comprehension) 260 (Zhong et al., 2023; Wang et al., 2021)
MedMCQA Valid MedMCQA Medicine 6,150 (Pal et al., 2022)
SAT-English AGI Eval English 206 (Zhong et al., 2023)
SAT-Math AGI Eval Math 220 (Zhong et al., 2023)
@@
Then, we converted the benchmark problems from their original data format into a standardized data structure using the JSON Lines (JSON-L) format (Ward, ). Our standardized set of exams allowed us to use the exams interchangeably without modifying the code in the test harness. See Appendix B - Figure 8 for a sample of an MCQA problem.
@@
Finally, we created two MCQA exams of different sizes. we created a large exam with 1,000 questions by randomly sampling 100 problems from each of the ten problem sets. This 1,000-question (large) exam was used with GPT-3.5 to perform our hypothesis test and a detailed analysis of temperature across problem domains and prompt types.
@@
However, we also created a smaller exam of 100 questions by randomly sampling ten questions from each of the ten domain-specific problem sets. This 100-question (small) exam was used for our high-level analysis of sampling temperature across all four models [^3] We used the smaller 100-question exam for GPT-4, Llama 7B and Llama 70B due to cost considerations..
2.2 Models
The LLM models used in this research project consist of four commonly used foundational LLMs. To complement our analysis, we also performed experiments using five prompts created with commonly used prompt-engineering techniques.
First, we reviewed the prior literature to identify candidate LLMs commonly used for problem-solving tasks. We limited our candidate models to those that allowed the model’s sampling temperature to be specified via their API. These LLMs included GPT-3.5, GPT-4, Llama 2 7B, and Llama 2 70B (OpenAI, 2022, 2023b, 2023c; Touvron et al., 2023).
As a result, we selected the following five prompt-engineering techniques to construct our system prompts:
  • Domain Expertise – the system prompt specifies that the LLM is an expert in the problem domain of the exam (e.g., “medicine”) or the topic of the problem (e.g., “anatomy”) (White et al., 2023).
  • Self-recitation – the system prompt instructs the LLM to recite its own internal knowledge about the problem before answering the question (Sun et al., 2023; White et al., 2023).
  • Chain-of-Thought (CoT) – the system prompt instructs the LLM to “think step-by-step” to encourage it to reason through the problem procedurally (Kojima et al., 2022; Wei et al., 2022).
  • Self-criticism – the system prompt instructs the LLM to critically evaluate its previous thoughts to identify any errors in facts, logic or reasoning (Huo et al., 2023; Wang et al., 2023b).
  • Composite – the system prompt combines domain expertise, self-recitation, chain-of-thought, and self-criticism.
Finally, we provided the LLM with a single example problem-and-solution pair for one-shot in-context learning. The example solution was adapted for each prompt based on the prompt-engineering technique used. For example, the chain-of-thought prompt included a chain-of-thought in its solution. See Appendix A - Figure 7 for a sample prompt.
@@ 2.3 Metrics
To test our hypothesis, we measured the LLM’s correct-answer accuracy as our primary performance metric. To further support our findings, we also measured the similarity of the LLM’s responses using a series of text-similarity metrics.
These metrics are defined as follows:
  • Accuracy – the ratio of correctly answered MCQA problems to the total number of problems attempted
  • Jaccard similarity – the ratio of the intersection to the union of word sets in the output text (Jaccard, 1912)
  • Bag-of-Words (BoW) similarity – comparison of the frequency of each word, ignoring order (Harris, 1954)
  • TF-IDF similarity – comparison of word frequency weighted by inverse document frequency (Jones, 1972)
  • Levenshtein similarity – the number of edits needed to change one text into the other (Levenshtein, 1966)
  • BLEU score – comparison of similarity based on n-gram overlap (Papineni et al., 2001)
  • SBERT similarity – semantic similarity computed using Sentence-BERT embeddings (Reimers & Gurevych, 2019)
@@ 2.4 Analysis
We performed an analysis of correct-answer accuracy across temperatures ranging from 0.0 to 1.0. in increments of 0.1 using GPT-3.5 with the CoT prompt and the 1,000-question exam. We then performed a Kruskal-Wallis test to evaluate the statistical significance of any changes in accuracy as a function of temperature [^4] We used Kruskal-Wallis because the data (i.e., correct-answer accuracy by question) were not normally distributed. Rather, they were bimodally distrbuted with centers at 0.0 and 1.0.. We evaluated our hypotheses at a significance level $\alpha=0.05$.
@@ 3 Results
Our analysis revealed that the mean accuracy remained relatively stable across all sampling temperatures.
@@
The Kruskal-Wallis test for GPT-3.5 on the 1,000-question exam yielded the following results:
  • Baseline: $H(10)=0.800,p=1.000$
  • Domain Expertise: $H(10)=2.508,p=0.991$
  • Self-Recitation: $H(10)=4.592,p=0.917$
  • Chain-of-Thought: $H(10)=10.407,p=0.406$
  • Composite: $H(10)=9.454,p=0.490$
Given a significance level of 0.05, the higher p-values indicate that the differences in accuracy across temperatures were not statistically significant. Thus, we did not find sufficient evidence to reject the null hypothesis.
Figure 1: Accuracy by temperature and prompt for GPT-3.5 with 1,000 questions. Performance remains relatively stable across all temperatures and prompts. However, there is a non-significant decrease in performance as a function of temperature.
In addition, we saw similar results for GPT-4 using the smaller 100-question exam. The two Llama models did not perform better than statistically random chance on the 100-question exam, so their results had to be excluded from our analysis [^5] The low accuracy of the Llama models was the result of it producing incorrectly formatted answers (7B: 38%; 70B: 57%) and correctly formatted but incorrect answers (7B: 57%; 70B: 19%).. See Figure 2.
Figure 2: Accuracy by temperature and model. Performance remains stable across sampling temperatures for all four LLMs on the 100-question MCQA exam. However, both Llama 2 models performed no better than statistically random guesses.
Figure 3: Accuracy of GPT-3.5 by temperature, prompt, and exam. Performance remains relatively stable across all temperatures regardless of the system prompt or the problem domain for GPT-3.5 on all 1,000 multiple-choice questions.
These results suggest that changes to sampling temperature in the range of 0.0 to 1.0 do not have a statistically significant effect on LLM performance in problem-solving. However, visual inspection of GPT-3.5 accuracy with 1,000 questions indicates a non-significant decrease in performance as a function of temperature – which matches our intuition about randomness in sampling temperature. See Figure 1.
Visual examination of the performance of GPT-3.5 across prompts and problem domains also supported these findings. We cannot identify any discernible trends in correct-answer accuracy for GPT-3.5 as a function of temperature when segmented by prompt type or exam. See Figure 3.
However, beyond a temperature of 1.0, problem-solving performance begins to decrease significantly starting at a temperature of 1.2 until the model produces statistically random guesses at a temperature of 1.4. These results also match our intuition of randomness in sampling temperature. See Figure 4.
Figure 4: Accuracy by temperature for GPT-3.5 with 100 questions beyond a temperature of 1.0. Performance begins to drop significantly at a temperature of 1.2 until performance becomes no better than statistically random guesses at 1.4.
@@
To further support our results, we analyzed text variability as a function of temperature. Our results show a clear trend of decreasing text similarity (thus increasing text variability) as temperature increases. This confirms that the sampling-temperature hyperparameter was behaving as expected.
Figure 5: Text similarity by temperature, prompt, and exam - TF-IDF text similarity decreases as temperature increases regardless of the prompt or problem domain. Note that the baseline and domain-expert prompts produce short LLM responses, so their decrease in text similarity across temperatures is minimal. However, math problems frequently cause the LLM to automatically produce a CoT.
Figure 6: Text similarity by temperature and metric. Text similarity decreases as temperature increases for GPT-3.5 on all 1,000 questions using all six text-similarity metrics.
These results are consistent with our understanding of sampling temperature, indicating that higher temperatures produce more widely varied outputs. In addition, these results hold regardless of the LLM model, prompt-engineering technique, or problem domain. See Figures 5 and 6.

### v2

@@ Abstract
In this research study, we empirically investigate the effect of sampling temperature on the performance of Large Language Models (LLMs) on various problem-solving tasks. We created a multiple-choice question-and-answer (MCQA) exam by randomly sampling problems from standard LLM benchmarks. Then, we used nine popular LLMs with five prompt-engineering techniques to solve the MCQA problems while increasing the sampling temperature from 0.0 to 1.6. Despite anecdotal reports to the contrary, our empirical results indicate that changes in temperature from 0.0 to 1.0 do not have a statistically significant impact on LLM performance for problem-solving tasks. In addition, these results appear to generalize across LLMs, prompt-engineering techniques, and problem domains. All code, data, and supplemental materials are available on GitHub at: https://github.com/matthewrenze/jhu-llm-temperature.
@@ 1.1 Background
In recent years, Large Language Models (LLMs) have revolutionized the field of artificial intelligence. The availability of open-source LLMs and pay-per-use APIs has allowed engineers to incorporate LLMs in their AI systems. However, prompt engineering and hyperparameter tuning are required to work effectively with LLMs.
Prompt-engineering techniques help LLMs solve complex problems, avoid hallucinations, and provide more accurate responses. For example, we can use techniques like chain-of-thought, tree-of-thought, self-criticism, and self-consistency to improve LLM performance [1, 2].
In addition, several inference hyperparameters can be adjusted to modify the LLM’s output at runtime. For example, hyperparameters like sampling temperature, top-k sampling, repetition penalty, and maximum token length all affect the LLMs output and performance [3, 4, 5].
Despite significant interest in LLMs and progress in LLM best practices, many open questions remain about optimal prompt-engineering techniques and inference hyperparameters for LLMs. To complicate matters, various local optima may exist for LLM, prompt type, and problem domain [5].
@@ 1.2 Sampling Temperature
Sampling temperature is a hyperparameter of an LLM used in a temperature-based sampling process. It controls the randomness of the model’s output at inference time [6, 7, 8, 5].
During each step of an LLM’s decoding process, the LLM uses the previous tokens to choose the next output token. The final layer of the LLM uses a softmax function to convert raw scores (logits) into probabilities.
@@
A lower temperature makes the output of the LLM more deterministic, thus favoring the most likely predictions. This conservativeness is captured by the model’s tendency to produce more repetitive, focused, and less diverse output based on the patterns most commonly seen in the training data [7, 8, 5].
A higher temperature increases the randomness of the output, thus favoring more “creative” predictions. This creativity is captured by the model’s willingness to explore more unconventional and less likely outputs. Higher temperatures can lead to novel text, diverse ideas, and creative solutions to problems [7, 8, 5].
@@ 1.3 Choosing a Sampling Temperature
Within the prompt-engineering community, there are a variety of opinions and best practices regarding the ideal sampling temperature for various problem-solving tasks [9, 10].
Low sampling temperatures are recommended for tasks requiring precision and factual accuracy, such as technical writing, code generation, or question-answering [11, 12]. However, higher temperatures are recommended for tasks requiring creativity, such as writing poetry, creating stories, or brainstorming.
Higher temperatures also increase the probability of model hallucination. Hallucination is a phenomenon where an LLM produces statistically probable responses that are factually incorrect or nonsensical. As a result, optimal temperature selection is also a balance between creativity and hallucination [13].
Practical guidelines for choosing a sampling temperature for a specific task or problem domain are often vague or anecdotal. Prompt engineering guides often provide hypothetical examples of optimal sample temperatures for various tasks. However, they rarely cite any sources or provide empirical evidence. [^1] A few empirical studies exist that indicate sampling temperature does have an effect on LLM performance on some types of problem-solving tasks (e.g., code generation, engineering exams, etc.) [11, 14, 12].
As a result, the current state of choosing the optimal sampling temperature for specific problems is largely based on guesswork, gut instinct, non-systematic experimentation, and iterative refinement. [^2] For example, OpenAI’s GPT-3.5 API allowed users to set the sampling temperature from 0.0 to 1.0 with a default of 0.7. GPT-4’s API expanded this range from 0.0 to 2.0 with a default of 1.0. No explanation from OpenAI has been provided for these default values or their change from GPT-3.5 to GPT-4 [15]. [^3] Even the GPT-4 Technical Report explains that the authors used their “best-guess” when choosing sampling temperatures while evaluating GPT-4 on various benchmarks. See Appendix A in the GPT-4 Technical Report [16].
@@ 2 Methods
2.1 Models
The LLM models used in this research project comprise nine widely-used foundational LLMs. To complement our analysis, we also conducted experiments using five prompts created using commonly used prompt-engineering techniques.
First, we reviewed the prior literature to identify candidate LLMs commonly used for problem-solving tasks. We limited our candidate models to those that allowed the model’s sampling temperature to be specified via their API [17, 18, 16, 4]. See Table 1 for a list of LLMs used in the experiment.
Name Vendor Released License Source
Claude 3 Opus Anthropic 2024-03-04 Closed [19, 20]
Command R+ Cohere 2024-04-04 Open [21, 22]
Gemini 1.0 Pro Google 2023-12-06 Closed [23, 24]
Gemini 1.5 Pro (Preview) Google 2024-02-15 Closed [25, 26]
GPT-3.5 Turbo OpenAI 2022-11-30 Closed [17, 27]
GPT-4 OpenAI 2023-03-14 Closed [18, 16]
Llama 2 7B Chat Meta 2023-07-18 Open [28, 4]
Llama 2 70B Chat Meta 2023-07-18 Open [28, 4]
Mistral Large Mistral AI 2024-02-26 Closed [29]
Table 1: LLMs used in the experiment.
As a result, we selected five prompt-engineering techniques to construct our system prompts:
  • Domain Expertise – the system prompt specifies that the LLM is an expert in the problem domain of the exam (e.g., “medicine”) or the topic of the problem (e.g., “anatomy”) [2].
  • Self-recitation – the system prompt instructs the LLM to recite its own internal knowledge about the problem before answering the question [30, 2].
  • Chain-of-Thought (CoT) – the system prompt instructs the LLM to “think step-by-step” to encourage it to reason through the problem procedurally [31, 32].
  • Composite – the system prompt combines domain expertise, self-recitation, chain-of-thought, and adds self-criticism [33, 34].
Finally, we provided the LLM with a single example problem-and-solution pair for one-shot in-context learning. The example solution was adapted for each prompt based on the prompt-engineering technique used. For example, the CoT prompt included a chain of thought in its solution. See Figure 10 in the Appendix for a sample prompt.
@@ 2.2 Data
The test dataset used in this research study consists of a series of Multiple-Choice Question-and-Answer (MCQA) exams derived from widely used LLM performance benchmarks.
@@
Next, we selected a set of problems that covered a range of problem domains (e.g., math, science, law, etc.) and difficulty levels (e.g., secondary school, university, etc.) These problem sets can be seen in Table 2.
Problem Set Benchmark Domain Questions License Source
ARC Challenge Test ARC Science 1,173 CC BY-SA [35]
AQUA-RAT AGI Eval Math 254 Apache v2.0 [36]
Hellaswag Val Hellaswag Common Sense Reasoning 10,042 MIT [37]
LogiQA (English) AGI Eval Logic 651 GitHub [36, 38]
LSAT-AR AGI Eval Law (Analytic Reasoning) 230 MIT [36, 39]
LSAT-LR AGI Eval Law (Logical Reasoning) 510 MIT [36, 39]
LSAT-RC AGI Eval Law (Reading Comprehension) 260 MIT [36, 39]
MedMCQA Valid MedMCQA Medicine 6,150 MIT [40]
SAT-English AGI Eval English 206 MIT [36]
SAT-Math AGI Eval Math 220 MIT [36]
@@
Table 2: Problem sets used to create the multi-domain MCQA exam.
Note: The GitHub repository for LogiQA does not include a license file. However, both the paper and readme.md file states that "The dataset is freely available."
@@
Then, we converted the benchmark problems from their original data format into a standardized data structure using the JSON Lines (JSON-L) format [41]. Our standardized set of exams allowed us to use the exams interchangeably without modifying the code in the test harness. See Figure 11 in the Appendix for a sample of an MCQA problem.
@@
Finally, we created two MCQA exams of different sizes. We created a large exam with 1,000 questions by randomly sampling 100 problems from each of the ten problem sets. This 1,000-question (large) exam was used with GPT-3.5 to perform a detailed analysis of temperature across problem domains.
@@
Additionally, we created a smaller exam of 100 questions by randomly sampling ten questions from each of the ten domain-specific problem sets. This 100-question (small) exam was used for our high-level analysis of sampling temperature across all nine models, all five prompt-engineering techniques, and extended temperature range (0.0-1.6). [^4] We used the smaller 100-question exam due to cost and runtime considerations.
2.3 Process
Our experiment was designed to test the problem-solving performance of LLMs across ten models, five prompt-engineering techniques, ten problem domains, 100 problems within each problem domain, and all viable sampling temperatures. For each combination of model, prompt, exam, and temperature, we instructed the LLM to answer each question ten times so we could assess the average correct-answer accuracy.
The full experiment setup can be seen in Figure 1 and Algorithm 1. However, due to cost and runtime considerations, we conducted a subset of the full experiment designed to capture the most valuable information as efficiently as possible.
Figure 1: Diagram of the full experiment.
Algorithm 1 Full LLM Temperature Experiment
1:for each model $m$ in $M$ do $\triangleright$ 10 models
2:   for each prompt $p$ in $P$ do $\triangleright$ 5 prompts
3:     for each exam $e$ in $E$ do $\triangleright$ 10 exams
4:       for each temperature $\tau$ in $T$ do $\triangleright$ 16 temps
5:         for each problem $q$ in $Q$ do $\triangleright$ 100 probs
6:            for each attempt $a$ in $A$ do $\triangleright$ 10 attempts
7:              Create the prompt
8:              Answer the question
9:              Record the answer
10:            end for
11:         end for
12:       end for
13:       Save the results
14:     end for
15:   end for
16:end for
17:Process the results
18:Analyze the results
First, we instructed GPT-3.5 to complete the 100-question (small) exam using the CoT prompt with temperatures ranging from 0.0 to 2.0 in increments of 0.1. This allowed us to determine the range of viable sampling temperatures to explore.
Performance began to drop rapidly after a temperature of 1.0 until the generated text became incoherent at 1.6. As a result, we stopped the initial temperature sweep at 1.6 and limited the rest of our sweeps from 0.0 to 1.0.
Next, we instructed the other eight LLMs to complete the 100-question (small) exam using the CoT prompt with temperatures from 0.0 to 1.0. This allowed us to determine if the results generalize to other LLMs.
Then, we instructed GPT-3.5 to complete the 100-question (small) exam using each of the five prompts over temperatures from 0.0 to 1.0. This allowed us to determine if the results generalize over various prompt-engineering techniques.
Finally, we instructed GPT-3.5 to complete the 1,000-question (large) exam using the CoT prompt with temperatures from 0.0 to 1.0. This allowed us to determine if the results were domain-specific or generalized across problem domains.
@@ 2.4 Metrics
To test our hypothesis, we measured the LLM’s correct-answer accuracy as our primary performance metric. For each combination of model, prompt, exam, and temperature, we calculated the accuracy as the number of correct answers from ten attempts at each problem. Then, we computed the average (mean) accuracy across all problems.
To further support our findings, we also measured the similarity of the LLM’s responses using a series of text-similarity metrics. These metrics are defined as follows:
  • Jaccard similarity – the ratio of the intersection to the union of word sets in the output text [42].
  • Bag-of-Words (BoW) similarity – comparison of the frequency of each word, ignoring order [43].
  • TF-IDF similarity – comparison of word frequency weighted by inverse document frequency [44].
  • Levenshtein similarity – the number of edits needed to change one string of text into the other [45].
  • BLEU score – comparison of similarity based on n-gram overlap [46].
  • SBERT similarity – semantic similarity computed using Sentence-BERT embeddings [47].
@@ 2.5 Analysis
We used the Kruskal-Wallis test to evaluate the statistical significance of any changes in accuracy as a function of temperature [48]. We chose the Kruskal-Wallis test because the data (i.e., correct-answer accuracy by question) were not normally distributed. Rather, they were bimodally distributed with centers at 0.0 and 1.0.
@@ 3 Results
Our analysis revealed that the problem-solving performance of LLMs remained relatively stable across sampling temperatures from 0.0 to 1.0 for all LLMs, prompt-engineering techniques, and problem domains. Using GPT-3.5 with a CoT prompt on the 1,000-question exam from 0.0 to 1.0, the Kruskal-Wallis test yielded $H(10)=10.439,p=0.403$.
First, we analyzed the performance of GPT-3.5 using the CoT prompt on the 100-question exam. Accuracy remained stable over temperatures from 0.0 to 1.0. However, after a temperature of 1.0, the text rapidly became incoherent, and the accuracy began to drop until it reached zero around a temperature of 1.6. See Figure 2.
Figure 2: Accuracy by temperature from 0.0 to 1.6 for GPT-3.5 using the CoT prompt on the 100-question exam.
Second, we analyzed the performance of all nine LLMs using the CoT prompt on the 100-question exam. Accuracy also remained stable across all of the LLMs, except for Llama 2 7B. The performance of most LLMs showed a gradual (non-significant) decrease in performance as a function of temperature. See Figure 3 and Table 3.
Llama 2 7B did not perform better than statistically random guesses. Its poor performance was due to generating incorrectly formatted answers (39%) and correctly formatted but incorrect answers (36%). Its all-or-nothing behavior at a temperature of 0.0 versus more random behavior at 0.1 to 1.0 led to a much lower, yet still non-significant, p-value.
Figure 3: Accuracy by temperature and model using the
CoT prompt on the 100-question exam.
Model H(10) p-value
Claude 3 Opus 1.735 0.998
Command R+ 1.771 0.998
Gemini Pro 1.0 7.379 0.689
Gemini Pro 1.5 3.119 0.978
GPT-3.5 Turbo 2.042 0.996
GPT-4 3.789 0.956
Llama 2 70B 3.677 0.961
Llama 2 7B 17.086 0.072
Mistral Large 3.069 0.980
@@
Table 3: Kruskal-Wallis test results by model using the CoT prompt on the 100-question exam.
Third, we analyzed the performance of GPT-3.5 using each of the five prompts on the 100-question exam. Accuracy remained stable for all temperatures across all prompt-engineering techniques. The CoT prompt outperformed the other four prompts. As a result, we used the CoT prompt for all single-prompt experiments. See Figure 4 and Table 4.
Figure 4: Accuracy by temperature and prompt for
GPT-3.5 on the 100-question exam.
Prompt H(10) p-value
Baseline 0.420 1.000
Domain Expert 0.548 1.000
Self-recitation 1.403 0.999
Chain of Thought 2.042 0.996
Composite 1.000 1.000
Table 4: Kruskal-Wallis test results by prompt for
GPT-3.5 on the 100-question exam.
Finally, we analyzed the performance of GPT-3.5 using the CoT prompt on all ten exams. Accuracy remained stable for all temperatures across all problem domains based on visual analysis. However, the LSAT-AR and SAT-Math exams showed statistically significant differences in the Kruskal-Wallis p-values. See Figure 5 and Table 5.
We performed the Dunn-Bonferroni test on the LSAT-AR and SAT-Math results [49]. It revealed that the all-or-nothing behavior of responses generated at a temperature of 0.0 versus the more random responses from 0.1 to 1.0 caused the anomaly. The correct-answer accuracy for each individual problem varied significantly when evaluated pair-wise across temperatures. However, the average accuracy for all problems remained similar across temperatures.
Figure 5: Accuracy by temperature and exam for GPT-3.5
using the CoT prompt.
Exam H(10) p-value
AQUA-RAT 10.320 0.413
ARC Challenge 16.390 0.089
Hellaswag 4.473 0.924
LogiQA 3.208 0.976
LSAT-AR 37.874 < 0.001
LSAT-LR 7.816 0.647
LSAT-RC 4.037 0.946
MedMCQA 2.334 0.993
SAT-English 3.937 0.950
SAT-Math 21.276 0.019
Table 5: Kruskal-Wallis test results by exam for GPT-3.5 using the CoT prompt.
@@
To further support our results, we analyzed text variability as a function of temperature. Our findings show a clear trend of decreasing text similarity (thus increasing text variability) as temperature increases. Text similarity decreases rapidly after a temperature of 1.0, corresponding to the rapid decrease in accuracy observed above $\tau=1.0$. See Figure 7.
These results are consistent with our understanding of sampling temperature, indicating that higher temperatures produce more widely varied outputs. Furthermore, these results hold regardless of the LLM, prompt-engineering technique, or problem domain. See Figures 7, 9, and 9.
Figure 6: Text similarity by temperature and metric for GPT-3.5 using CoT prompting on the 100-question exam over sampling temperatures from 0.0 to 1.6.
Figure 7: TF-IDF text similarity by temperature and model using the CoT prompt on the 100-question exam over sampling temperatures from 0.0 to 1.0
Figure 8: TF-IDF text similarity by temperature and prompt for GPT-3.5 on the 100-question exam over sampling temperatures from 0.0 to 1.0.
Figure 9: TF-IDF text similarity by temperature and exam for GPT-3.5 using the CoT prompt over sampling temperatures from 0.0 to 1.0